OpenStudy (anonymous):

factor completely x^(17/3)-25x^(11/3)

OpenStudy (anonymous):

is it x^(11/3)(x-5)(x+5)

hero (hero):

You should learn how to use the equation editior. I'm having trouble interpreting your math

OpenStudy (anonymous):

basically the problem is x to the 17/3 power minus 25x to the 11/3 power

hero (hero):

bizzarro

hero (hero):

Doesn't this involve logs?

hero (hero):

According to wolfram, your answer is correct. I'm still scratching my head though

hero (hero):

How did you get your answer?

OpenStudy (anonymous):

the x to (11/3) power im not sure thats the problem lol

OpenStudy (anonymous):

how do you use wolframalpha?

hero (hero):

I think I know how it is done

hero (hero):

apparently you factor out x^11/3 leaving you with x^(11/3)(x^2 - 25) then that becomes x^(11/3)(x + 5)(x - 5)

hero (hero):

somehow when you factor out the x^(11/3) you end up with (x^(6/3) - 25) which becomes (x^2 - 25)

hero (hero):

I guess because 17/3 - 11/3 = 6/3

hero (hero):

To use wolfram alpha, write: solve(x^(17/3)-25x^(11/3)=0,x)

hero (hero):

But even the steps for that is kind of shoddy. I don't like it

OpenStudy (anonymous):

okay thanks very much